# Burnside type results for fusion categories

### v10

@@ 4.1. Inequalities for $P$ and ${\widehat{P}}$
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.23. Then $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|\leq 1$. Moreover one has equality if and only if $x_{i}\in G{(H,\mathcal{B})}$.
@@
First suppose that ${(H,\mathcal{B})}$ is RN. Then $|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$ by Frobenius-Perron theory. This implies that $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|={\prod_{j=1}^{m}}|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$. Moreover one has equality if and only if $|{\frac{x_{i}}{d_{i}}}(\mu_{j})|=1$ for all $j\in\mathcal{J}$. Then Lemma 2.20 implies that $x_{i}$ is a grouplike element.
Now suppose that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then since $\mu_{j}=\sum_{i=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}}){\widehat{E}}_{i}$ it follows that $|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$ since $\mathrm{FPdim}(\mu_{j})=1$. This implies again that $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|\leq 1$. Moreover one has equality if and only if $|{\frac{x_{i}}{d_{i}}}(\mu_{j})|=1$ for all $j\in\mathcal{J}$. Then, again, Lemma 2.20 implies that $x_{i}$ is a grouplike element. ∎
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. One has that ${\widehat{P}}^{2}$ is an idempotent if and only if the non-vanishing elements $x_{i}$ of ${(H,\mathcal{B})}$ satisfy ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$.
@@
Note that Equation (4.1) one has ${\widehat{P}}^{4}={\widehat{P}}^{2}$ if and only if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})^{4}={\widehat{P}}({\frac{x_{i}}{d_{i}}})^{2}$ for all $i\in\mathcal{I}$. Therefore if ${\widehat{P}}^{2}$ is idempotent then either ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=0$ or ${\widehat{P}}({\frac{x_{i}}{d_{i}}})^{2}=1$ for all $i\in\mathcal{I}$. In the first case, since ${\widehat{P}}({\frac{x_{i}}{d_{i}}})={\prod_{j=1}^{m}}\mu_{j}({\frac{x_{i}}{d_{i}}})$ it follows that $x_{i}$ is a vanishing element. On the other hand if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})^{2}=1$ implies ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$. ∎
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.23. Then ${\widehat{P}}^{2}$ is idempotent if and only if ${(H,\mathcal{B})}$ is Burnside.
@@
In this case, by Proposition 4.11, one has $|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$ for all $j\in\mathcal{J}$ Thus ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$ implies $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|=1$ and Lemma 2.20 implies that $x_{i}\in G{(H,\mathcal{B})}$.
Conversely, suppose that ${(H,\mathcal{B})}$ is Burnside. It follows from Equation (4.1) that ${\widehat{P}}=\sum_{x_{i}\in G{(H,\mathcal{B})}}{\widehat{P}}({\frac{x_{i}}{d_{i}}}){\widehat{E}}_{i}$. Note that ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\det(L_{\frac{x_{i}}{d_{i}}})$ the operator of left multiplication by ${\frac{x_{i}}{d_{i}}}$ on $H$. By Corollary 2.32 this is the determinant of a permutation matrix and therefore equal to ${\mathrm{sgn}}(x_{i})=\pm 1$. ∎
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Define also $P={\prod_{i=1}^{m}}{\frac{x_{i}}{d_{i}}}$. One has
(4.10) $P=\sum_{j=1}^{m}\mu_{j}(P)F_{j}.$
The dual version of the last three results imply the following:
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.23. Then $|\mu_{j}(P)|\leq 1$ and one has equality if and only if $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$.
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. One has that $P^{2}$ is an idempotent if and only if the non-vanishing elements $\mu_{j}\in{\widehat{\mathcal{B}}}$ satisfy $\mu_{j}(P)=\pm 1$.
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.23. One has that $P^{2}$ is an idempotent if and only if ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside.
@@ 4.2. The analogue of Burnside’s theorem
We prove the following more general version of Theorem 1.2. Its proof follows the same lines as the proof of [Bur23, Theorem 2].
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian rational normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is ${h}$-integral. Let $x_{i}\in\mathcal{B}$ be an element such that $h_{i}|d_{i}|^{2}>1$ and
@@
for all $j\in\mathcal{J}$. Then there is $\mu_{j}\in{\widehat{\mathcal{B}}}$ such that $\mu_{j}(x_{i})=0$.
@@
For any $1\leq i\leq m$, denote $\mathcal{T}_{i}:=\mathcal{T}(x_{i})=\{j\in\mathcal{J}\;|\;\mu_{j}(x_{i})=0\}$ and ${\mathcal{D}}_{i}:=\mathcal{J}\setminus(\mathcal{T}_{i}\cup\{1\})$. One needs to show that $\mathcal{T}_{i}\neq\emptyset$, for any basis element $x_{i}\in\mathcal{B}$ satisfying the hypothesis from the theorem.
Note that the second orthogonality relation can be written as:
(4.12) $\sum_{j=1}^{m}\widehat{h}_{j}|\mu_{j}(x_{i})|^{2}=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}}$
Since $\widehat{h}_{1}=1$, for $j=1$, the above equation can be rewritten as:
@@
On the other hand, note that
$n{(H,\mathcal{B},\mu_{1})}=\sum_{j=1}^{m}\widehat{h}_{j}=1+\sum_{j\in\mathcal{T}_{i}}\widehat{h}_{j}+\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j}.$
Therefore Equation (4.13) can be written as:
@@
since then, it follows that $\frac{1+\sum\limits_{j\in\mathcal{T}_{i}}\widehat{h}_{j}}{h_{i}|d_{i}|^{2}}\geq 1$, i.e. $1+\sum\limits_{j\in\mathcal{T}_{i}}\widehat{h}_{j}\geq h_{i}|d_{i}|^{2}$. Since $h_{i}|d_{i}|^{2}>1$, it follows that $\mathcal{T}_{i}\neq\emptyset$.
The inequality from Equation (4.15) can be written as
(4.16) $\frac{1}{\sum\limits_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j}}(\sum_{j\in{\mathcal{D}}_{i}}{\widehat{h}_{j}h_{i}|\mu_{j}(x_{i})|^{2}})\geq 1.$
@@
(4.17) $\frac{1}{\sum\limits_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j}}(\sum_{j\in{\mathcal{D}}_{i}}{\widehat{h}_{j}h_{i}|\mu_{j}(x_{i})|^{2}})\geq\prod_{j\in{\mathcal{D}}_{i}}\bigg{(}h_{i}\big{(}|\mu_{j}(x_{i})|^{2}\big{)}^{\widehat{h}_{j}}\bigg{)}^{\frac{1}{|{\mathcal{D}}_{i}|}},$
where $|{\mathcal{D}}_{i}|:=\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j}$. Note that Equation (3.1) implies that the set ${\mathcal{D}}_{i}$ is stable under the Galois group $\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$. This in turn implies that the product
@@
is fixed by the Galois group $\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$, since ${\widehat{h}_{\tau(j)}}=\widehat{h}_{j}$ by Lemma 3.8 . It follows that $P_{i}$ is a rational number. On the other hand, each factor of $P_{i}$ is an algebraic integer (since $\widehat{h}_{j}\in\mathbb{Z}_{{}_{>0}}$), and therefore, the entire product is an integer. Since it is also positive, it must be at least $1$. ∎
Note that Theorem 1.2 follows from the above theorem, since in the case of a fusion ring, $h_{i}=1$ and $\mu_{j}(x_{i})\in\mathbb{A}$, for all $(i,j)\in{\mathcal{I}}\times{\mathcal{J}}$.
@@
Let ${(H,\mathcal{B})}$ be a rational ARN-hypergroup such that $\widehat{(H,\mathcal{B})}$ is ${h}$-integral. Let $x_{i}\in\mathcal{B}$ be an invertible element in $H$ satisfying also Equation (4.11), for all $1\leq j\leq m$. Then $x_{i}\in{G{(H,\mathcal{B})}}$ is a grouplike element.
@@
Note that ${(H,\mathcal{B})}$ satisfies the hypothesis of Theorem 4.17. Also, by hypothesis, Equation (4.11) holds for all $1\leq j\leq m$. Suppose that $x_{i}$ is not a grouplike element, i.e. $h_{i}|d_{i}|^{2}>1$, by Lemma 2.17. In this case, by Theorem 4.17, there is $\mu_{j}\in\widehat{B}$ such that $\mu_{j}(x_{i})=0$. Thus, if $x_{i}y=1$ then $10=\mu_{j}(x_{i})\mu_{j}(y)=$ 1, which is a contradiction. ∎
@@
Any abelian rational normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is ${h}$-integral and Equation (4.11) is satisfied for all $1\leq i,j\leq m$ is Burnside.
@@
By Proposition 4.3, it is enough to show that $\mathcal{B}_{1}={G{(H,\mathcal{B})}}$. Clearly, ${G{(H,\mathcal{B})}}\subseteq\mathcal{B}_{1}$. Suppose that $x_{i}\in\mathcal{B}$ is an element which is invertible in $H$. Since $x_{i}$ is not a vanishing element, it follows from Proposition 4.18 that $x_{i}\in{G{(H,\mathcal{B})}}$. ∎
@@
Let ${(H,\mathcal{B})}$ be a commutative fusion ring such that $\widehat{(H,\mathcal{B})}$ is ${h}$-integral. Then ${(H,\mathcal{B})}$ is Burnside.
@@
Any commutative Grothendieck ring of any weakly-integral fusion category is Burnside.
@@
It follows from above corollary by Remark 3.7. ∎
Now one can write the corresponding statement of Theorem 4.17 applied to the dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ of an abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$. We obtain the following:
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable ${h}$-integral hypergroup such that the dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is a rational hypergroup. Let $\mu_{j}\in{\widehat{\mathcal{B}}}$ be an element of ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ with $\widehat{h}_{j}>1$, and such that
(4.18) $\frac{\widehat{h}_{j}}{d_{i}^{2}}|\mu_{j}(x_{i})|^{2}\in\mathbb{A},$
for all $i\in\mathcal{I}$. Then, there is $x_{i}\in\mathcal{B}$ such that $\mu_{j}(x_{i})=0$.
@@
Note that the condition from Equation (4.11) applied on ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ becomes exactly Equation (4.18). ∎

### v11

@@ 4.1. Inequalities for $P$ and $^{\widehat{P}}$
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.30. Then $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|\leq 1$. Moreover, the equality holds if and only if $x_{i}\in G{(H,\mathcal{B})}$.
@@
First, suppose that ${(H,\mathcal{B})}$ is RN. Then $|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$ by Frobenius-Perron theory. This implies that $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|=\prod_{j\in\mathcal{I}}|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$. Moreover, the equality holds if and only if $|{\frac{x_{i}}{d_{i}}}(\mu_{j})|=1$ for all $j\in\mathcal{I}$, if and only if $x_{i}$ is grouplike, by Lemma 2.27.
Next, suppose that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Since $\mu_{j}=\sum_{i\in\mathcal{I}}\mu_{j}({\frac{x_{i}}{d_{i}}})\widehat{F}_{i}$, then $\{\mu_{j}({\frac{x_{i}}{d_{i}}})\}_{i\in\mathcal{I}}$ are the eigenvalues of $L_{\mu_{j}}$. By Frobenius-Perron theory, $|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq\mathrm{FPdim}(\mu_{j})=1$, as $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is normalized. Thus, $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|\leq 1$, again. The rest is as before. ∎
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then, ${\widehat{P}}^{2}$ is an idempotent if and only if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$, for all $x_{i}\in\mathcal{B}_{1}$.
@@
By Equation (4.1), ${\widehat{P}}^{4}={\widehat{P}}^{2}$ if and only if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})^{4}={\widehat{P}}({\frac{x_{i}}{d_{i}}})^{2}$ for all $i\in\mathcal{I}$. Therefore, ${\widehat{P}}^{2}$ is idempotent if and only if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=0$ or ${\widehat{P}}({\frac{x_{i}}{d_{i}}})^{2}=1$, for all $i\in\mathcal{I}$. Since ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\prod_{j\in\mathcal{I}}\mu_{j}({\frac{x_{i}}{d_{i}}})$, the second case occurs if and only if $x_{i}\in\mathcal{B}_{1}$, if and only if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$. ∎
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.30. Then, ${\widehat{P}}^{2}$ is an idempotent if and only if ${(H,\mathcal{B},\mu_{1})}$ is Burnside.
@@
By Proposition 4.12, ${\widehat{P}}^{2}$ is an idempotent if and only if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$, for all $x_{i}\in\mathcal{B}_{1}$. But, by assumption and Proposition 4.11, $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|=1$ if and only if $x_{i}\in G{(H,\mathcal{B})}$. Thus, ${\widehat{P}}^{2}$ idempotent implies that ${\mathcal{B}_{1}}\subset G{(H,\mathcal{B})}$. But $G{(H,\mathcal{B})}\subset\mathcal{B}_{1}$, so the equality holds, i.e. ${(H,\mathcal{B})}$ is Burnside.
Conversely, if ${(H,\mathcal{B})}$ is Burnside, then ${\widehat{P}}^{2}=\sum_{x_{i}\in G{(H,\mathcal{B})}}\widehat{F}_{i}$, by Corollary 4.5, which is an idempotent. ∎
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup, and $P:=\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}$. Then
(4.10) $P=\sum_{j\in\mathcal{I}}\mu_{j}(P)F_{j}.$
Here are the dual version of the last three results:
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.30. Then $|\mu_{j}(P)|\leq 1$. Moreover, the equality holds if and only if $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$.
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then, $P^{2}$ is an idempotent if and only if $\mu_{j}(P)=\pm 1$, for all $\mu_{j}\in\widehat{\mathcal{B}}_{1}$.
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.30. Then, $P^{2}$ is an idempotent if and only if ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside.
@@ 4.2. The analogue of Burnside’s theorem
We will prove the following generalization of Theorem 1.2, using the same approach as in [Bur23, Theorem 2].
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian rational normalizable hypergroup with ${h}$-integral dual. Let $x_{i}\in\mathcal{B}$. If $h_{i}|d_{i}|^{2}>1$, and
@@
for all $j\in\mathcal{I}$. Then $x_{i}\in\mathcal{B}_{0}$
@@
For any $i\in\mathcal{I}$, denote $\mathcal{T}_{i}:=\{j\in\mathcal{I}\;|\;\mu_{j}(x_{i})=0\}$ and ${\mathcal{D}}_{i}:=\mathcal{I}\setminus(\mathcal{T}_{i}\cup\{1\})$. We need to show that $\mathcal{T}_{i}\neq\emptyset$, for any $x_{i}\in\mathcal{B}$ satisfying the theorem’s hypothesis.
The second orthogonality relation (2.6) and Equation (2.11) implies
(4.12) $\sum_{j\in\mathcal{I}}\widehat{h}_{j}|\mu_{j}(x_{i})|^{2}=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}}.$
Since $\widehat{h}_{1}=1$, the above equation can be written as:
@@
On the other hand,
$n{(H,\mathcal{B},\mu_{1})}=\sum_{j\in\mathcal{I}}\widehat{h}_{j}=1+\sum_{j\in\mathcal{T}_{i}}\widehat{h}_{j}+\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j}.$
Therefore, Equation (4.13) can be written as
@@
since then, $\frac{1+\sum_{j\in\mathcal{T}_{i}}\widehat{h}_{j}}{h_{i}|d_{i}|^{2}}\geq 1$, i.e. $1+\sum\limits_{j\in\mathcal{T}_{i}}\widehat{h}_{j}\geq h_{i}|d_{i}|^{2}$; and $h_{i}|d_{i}|^{2}>1$ implies $\mathcal{T}_{i}\neq\emptyset$.
The inequality (4.15) can be written as
(4.16) $(\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j})^{-1}(\sum_{j\in{\mathcal{D}}_{i}}{\widehat{h}_{j}h_{i}|\mu_{j}(x_{i})|^{2}})\geq 1.$
@@
(4.17) $(\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j})^{-1}(\sum_{j\in{\mathcal{D}}_{i}}{\widehat{h}_{j}h_{i}|\mu_{j}(x_{i})|^{2}})\geq\bigg{(}\prod_{j\in{\mathcal{D}}_{i}}\big{(}h_{i}|\mu_{j}(x_{i})|^{2}\big{)}^{\widehat{h}_{j}}\bigg{)}^{(\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j})^{-1}}.$
By rational assumption, $h_{i}$ is rational and $\mu_{j}(x_{i})$ is an algebraic number. Equation (3.1) implies that the set ${\mathcal{D}}_{i}$ is stable under the action of $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$. It follows that the product
@@
is fixed by the action of $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, since ${\widehat{h}_{\tau(j)}}=\widehat{h}_{j}$ by Lemma 3.8. Thus $P_{i}$ is a rational number. On the other hand, each factor of $P_{i}$ is an algebraic integer (since $\widehat{h}_{j}\in\mathbb{Z}_{{}_{>0}}$), and therefore, the entire product is an integer. Since it is also positive, it must be at least 111. ∎
Note that Theorem 1.2 follows from the above theorem, since in the case of a fusion ring, $h_{i}=1$, and $\mu_{j}(x_{i})\in\mathbb{A}$, for all $i,j\in{\mathcal{I}}$.
@@
Let ${(H,\mathcal{B})}$ be a rational ARN-hypergroup with $h$-integral dual. Let $x_{i}\in\mathcal{B}$ be an invertible element in $H$ satisfying also Equation (4.11), for all $j\in\mathcal{I}$. Then $x_{i}\in{G{(H,\mathcal{B})}}$.
@@
Suppose that $x_{i}$ is not a grouplike element, i.e. $h_{i}|d_{i}|^{2}>1$, by Lemma 2.24. Then, all the assumptions of Theorem 4.17 are satisfied, so that, $x_{i}\in\mathcal{B}_{0}$, contradiction with $x_{i}$ invertible.∎
@@
Any rational ARN-hypergroup ${(H,\mathcal{B})}$ with $h$-integral dual, such that Equation (4.11) holds for all $i,j\in\mathcal{I}$, is Burnside.
@@
By Proposition 4.3, it is enough to show that ${\mathcal{B}_{1}}\subseteq{G{(H,\mathcal{B})}}$. Let $x_{i}\in\mathcal{B}_{1}$, thus $x_{i}$ is invertible, so by Proposition 4.18, $x_{i}\in{G{(H,\mathcal{B})}}$. ∎
@@
Any commutative fusion ring with $h$-integral dual is Burnside.
@@
A weakly-integral fusion category with a commutative Grothendieck ring is Burnside.
@@
It follows from Corollary 4.20 by Remark 3.7. ∎
We can write a kind of dual version of Theorem 4.17:
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable $h$-integral hypergroup with a rational dual. Let $\mu_{j}\in{\widehat{\mathcal{B}}}$ such that $\widehat{h}_{j}>1$ and
(4.18) $\widehat{h}_{j}\left|\mu_{j}(\frac{x_{i}}{d_{i}})\right|^{2}\in\mathbb{A},$
for all $i\in\mathcal{I}$. Then, $\mathcal{B}_{0}$ is non-empty.
@@
The condition from Equation (4.11) applied on $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ becomes exactly Equation (4.18). So we can apply Theorem 4.17 to $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$. Thus $\mu_{j}\in\widehat{\mathcal{B}}_{0}$, meaning the existence of $x_{i}\in\mathcal{B}$ such that $\mu_{j}(\frac{x_{i}}{d_{i}})=0$, so $x_{i}\in\mathcal{B}_{0}$. ∎